OpenStudy (anonymous):

y=x-2 y=-2x+7 Because both equations are equal to y, you can use the substation method to solve for x. Rewrite the system of equations so that the expressions containing x are equal to each other. ______________=_______________

OpenStudy (anonymous):

I got x-2=-2x+7 but how do you solve for x

OpenStudy (anonymous):

\(\Large x-2=-2x+7\\\Large x-2+2x=-2x+7+2x\\\Large 3x-2=7\\\Large 3x-2+2=7+2\\\Large 3x=9\\\Large x=\frac{9}{3}=2\)

OpenStudy (anonymous):

thank you sooo much!!

OpenStudy (anonymous):

ur welcome ^^